Solve: logbase2(1-x)+logbase2(4x+1)=0
Answer by stanbon(75887)

Solve: logbase2(1-x)+logbase2(4x+1) = 0
log2[(1-x)(4x+1)] = 0
(1-x)(4x+1) = 2^0
-4x^2+3x+1 = 1
4x^2-3x = 0
x(4x-3)=0
x = 0 or x = 3/4
Checking:
If x = 0:
log2(1)+log2)1) = 0
True
--------------
If x = 3/4
log2(1-(3/4)) + log2(4(3/4)+1) = 0
log2(1/4) + log2(4) = 0
-2 + 2 = 0
True
===========
Cheers,
Stan H.